# FedBaF: Federated Learning Aggregation Biased by a Foundation Model

### v1

@@ 4.1 General Convergence Analysis
Using the same restrictions on local and global learning rates placed by the FedAvg convergence analysis in [33], the aforementioned bounded gradient variance, bounded gradient dissimilarity, and L-smoothness assumptions ensure that our method converges to $\mathbf{w}^{*}$ faster than FedAvg. Similar convergence rate arguments for other FL methods with appropriately modified aggregation can be shown, as discussed in Sec. 3.
@@ 4.2 Effectiveness of FedBaF with Diverse Client Data
This section shows the impact of integrating a foundation model close to the optimal weights on the learning process and convergence behavior in non-IID environments.
@@
Aggregating the updated models from clients according to Alg. 1 Lines 8-10 forms the global model $\mathbf{w}^{\prime}_{t}$, and then,
$\delta_{t}\geq\|\mathbf{w}^{\prime}_{t}-\mathbf{w}^{*}\|.$
@@
Let $\mathbf{w}^{*}$ be a (bounded) local minimum of the global objective function in (1). Consider an FL algorithm that converges to $\mathbf{w}^{*}$ and let $\mathbf{w}^{\prime}_{t}$ be its global model. Consider FedBaF based on the same FL algorithm (with appropriately modified client updates and Lines 8-10 in Alg. 1) and let $\mathbf{w}_{t}$ be the FedBaF global model. FedBaF’s global model error has an upper bound of $\|\mathbf{w}_{t}-\mathbf{w}^{*}_{t}\|\leq\frac{\delta_{t}+\alpha_{t}\tau_{t}\gamma}{1+\alpha_{t}\tau_{t}}<\delta_{t}$.
@@ 5 Experimental Evaluations
In this section, we conduct a detailed evaluation of FedBaF’s performance on both local and global test datasets, comparing its performance to the no foundation model cases and weight initialization algorithms for training FL models.
@@
More detailed testing results are provided in Appendix C, including the 1) computational efficiency of FedBaF and 2) additional evaluations using foundation models of varying quality trained on different amounts of data and real-world foundation model weights that are officially available online.
@@ 5.1 Experimental Setup
Our experiments with popular image classification tasks encompass experiments using the CIFAR-10 and Rome Weather Image [34] datasets. We use Pre-ResNet and Vision Transformer [35] architectures as Vision Transformers are popular foundation model architectures known for achieving remarkable performance [36] for ImageNet challenges [35]. We train the foundation models on the Tiny ImageNet-200 and Weather Image [37] datasets. Our evaluations consider both IID and non-IID settings (see Appendix A[38, 39]. To demonstrate FedBaF’s generalizability to other tasks, we also evaluate it on a next-word prediction task using a Transformer language model pre-trained on the WikiText-2 dataset and tested on the Penn Treebank dataset.
@@
Details of other deep neural network architectures employed in our experiments and additional training specifics are provided in Tables 2 and 1 in Appendix A.
@@ 5.2 Experimental Results
Figures 3, 4, and 5 display the extensive test accuracy/perplexity evaluation results for FedBaF and our two baselines (no foundation model and weight initialization). We evaluate these three methods using both FedAvg and FedProx [31] as the base FL training algorithms. We incorporate FedBaF into FedProx by modifying Line 8 and the ClientUpdate routine in Algorithm 1, including both IID and non-IID settings, with one non-adversarial scenario and four adversarial scenarios. We use FedProx’s usual aggregation equation $\mathbf{w}^{\prime}_{t+1}\leftarrow\sum_{k\in S_{t}}\frac{n_{k}}{\sum_{k\in S_{t}}n_{k}}\mathbf{w}_{t+1}^{k}-\mu(\mathbf{w}_{t+1}^{k}-\mathbf{w}_{t})$. Here, $\mu$ represents the regularization term that controls the trade-off between the local and global objectives. FedBaF then incorporates the foundation model weights as in Lines 9-11 of Algorithm 1.
5.2.1 Testing Performance Enhancements
@@
Figures 3 and 4 respectively show the test accuracies of all three methods using FedAvg and FedProx for both Pre-ResNet and the Vision Transformer. In comparison to Pre-ResNet trained with no foundation model, FedBaF improved the global model’s accuracy by 1.3% for FedAvg and 1.6% for FedProx in IID scenarios, and by 21.8% for FedAvg and 22.6% for FedProx in non-IID scenarios. For the Vision Transformer, global performance improvements from FedBaF relative to no foundation model by 10.8% for FedAvg and 0.0% for FedProx in IID settings and 37.5% for FedAvg and 15.8% for FedProx in non-IID settings. We thus observe that both FedAvg and FedProx benefit from FedBaF’s inclusion of the foundation model, with particular benefits in the more challenging scenario with non-IID client data. Intuitively, non-IID data tends to slow FL convergence, which is mitigated by incorporating the foundation model.
@@
In the next-word prediction task using a Transformer, FedBaF significantly outperformed training with no foundation model, reducing perplexity by 76.0% with FedAvg, whereas weight initialization yielded a 67.8% decrease relative to training without a foundation model with FedAvg. Collectively, these findings indicate negligible differences between the test accuracies attained by FedBaF and those achieved with weight initialization, despite the fact that unlike weight initialization, FedBaF does not reveal the foundation model weights to FL clients. Moreover, FedBaF showed better testing performance than the weight initialization or no foundation model methods in 10 out of the 14 experiment settings.
5.2.2 Robustness to Attacks
@@
FedBaF remains effective in maintaining robustness against misclassification attacks, showing a more modest performance decline than both baseline methods.
@@
We next evaluate the robustness of FedBaF and our two baselines to adversarial clients. As the proportion of attacking clients ($\zeta$ in Figures 34, and 5) increases, test accuracy declines in all cases, as we would intuitively expect since even FedBaF is not designed to perfectly defend against these attacks, which become more effective as more clients act as attackers.
@@
The test accuracy for the no foundation model method especially drops significantly when 50% of the clients are attackers and the attack intensity $\lambda=5$. For Pre-ResNet under such attacks, there is a global performance decrease compared to the case with no attackers of 11.8% for FedAvg and 11.2% for FedProx in IID scenarios and 38.2% for FedAvg and 37.4% for FedProx in non-IID scenarios. For Vision Transformer, these accuracy drops are 32.4% for FedAvg and 34.2% for FedProx in IID and 21.9% for FedAvg and 25.0% for FedProx in non-IID scenarios. Thus, the no foundation model training method is vulnerable to attacks in both IID and non-IID settings, as we would expect since it has no built-in defenses.
@@
In contrast, FedBaF experiences a much more modest performance decline under attack, demonstrating its robustness. For Pre-ResNet, FedBaF’s global performance decreases by 6.0% for FedAvg and 6.4% for FedProx in IID settings and by 16.5% for FedAvg and 16.3% for FedProx in non-IID settings, where the decrease is again measured for the most intense attack ($\zeta=50\%,\lambda=5$ relative to no attack. These accuracy drops are less than half of those experienced by the no foundation model method. For Vision Transformer, FedBaF’s global performance decreases by only 4.9% for FedAvg and FedProx in IID settings and by 25% for both FedAvg and FedProx in non-IID settings.
@@
We finally see from the results for the weight initialization method that naïvely incorporating the foundation model still confers some robustness to attacks, though weight initialization reveals the foundation model weights to clients. Weight initialization shows a performance decrease of 6.4% for FedAvg and 6.9% for FedProx in IID settings and by 17.5% for FedAvg and 17.2% for FedProx in non-IID settings globally for Pre-ResNet. For Vision Transformer, the decreases are 14.6% for both FedAvg and FedProx in IID settings and by 18.4% for both FedAvg and FedProx in non-IID settings globally.
@@
These findings demonstrate that, similar to weight initialization, FedBaF offers considerable attack robustness compared to training with no foundation model. In 8 out of 12 cases, FedBaF shows the minimum decrease in performance across all three methods, indicating its superior effectiveness in maintaining robustness under adversarial conditions.
@@ 6 Conclusion
This section includes supplementary experimental results, analyzing the effect of varying foundation model quality on FedBaF’s performance and comparisons using the official pre-trained foundation model. Computational complexity is also compared to scenarios without foundation models and weight initialization.
@@ C.2 Using the Official Pre-Trained Foundation Model
C.3 Computational Complexity
@@
Compared to the no weight initialization cases, FedBaF requires significantly fewer computations across both IID and non-IID scenarios. Specifically, for Pre-ResNet IID and non-IID cases, computations are reduced by 88.1-91.4% and 89.7-92.4% for FedAvg, and by 88.2-91.4% and 89.8-92.4% for FedProx, respectively. Similarly, for Vision Transformer, IID and non-IID scenarios see a reduction of 96.9-97.3% and 93.9-96.8% for FedAvg, and 96.5-97.0% and 94.5-97.1% for FedProx in computations. Therefore, these findings indicate that FedBaF’s computational demands are relatively minimal despite the foundation model being integrated into every training round.
@@ F.2 Proof of Proposition 2
where we set $\gamma=\|\mathbf{w}_{\text{pre}}-\mathbf{w}^{*}\|$.
Since non-IID data can cause significant variance in local updates, we compare the derived bound to FedAvg, where the bound on the distance between $\mathbf{w}_{t}$ and $\mathbf{w}^{*}$ is $\delta_{t}$. By assumption, $\gamma\leq\delta_{t}$ for earlier rounds (small $t$). We get $\frac{\delta_{t}+\alpha_{t}\tau_{t}\gamma}{1+\alpha_{t}\tau_{t}}\leq\delta_{t}$, which is equivalent to
@@ G.1 Possibility of Extracting a Foundation Model
$\|\mathbf{w}_{t}-\mathbf{w}_{t}^{\prime}\|$ $=\left\|\frac{1}{1+\alpha\tau_{t}}(\mathbf{w}_{t}^{\prime}+\alpha\tau_{t}\mathbf{w}_{pre})-\mathbf{w}_{t}^{\prime}\right\|$
$=\frac{\alpha\tau_{t}}{1+\alpha\tau_{t}}\|\mathbf{w}_{t}^{\prime}-\mathbf{w}_{pre}\|$
$\|\mathbf{w}_{t}-\mathbf{w}_{pre}\|$ $=\left\|\frac{1}{1+\alpha\tau_{t}}\mathbf{w}_{t}^{\prime}+\alpha\tau_{t}\mathbf{w}_{pre}-\mathbf{w}_{pre}\right\|$
$=\frac{1}{1+\alpha\tau_{t}}\|\mathbf{w}_{t}^{\prime}-\mathbf{w}_{pre}\|$
At the onset of training, both distances are equivalent since we make a strategic choice for the weight $\alpha_{t}\tau_{0}$ to be approximately 2. This simplifies the initial update rule for the global model such that the initial global model weights, $\mathbf{w}_{0}$, are an unweighted average of the client’s updated model weights $\mathbf{w}_{0}^{\prime}$ and the foundation model weights $\mathbf{w}_{pre}$. As the training progresses, $\alpha\tau_{t}$ typically decays to less than 1. We deduce for $t>0$:
$\frac{\alpha\tau_{t}}{1+\alpha\tau_{t}}<\frac{1}{1+\alpha\tau_{t}}\implies\|\mathbf{w}_{t}-\mathbf{w}_{t}^{\prime}\|<\|\mathbf{w}_{t}-\mathbf{w}_{pre}\|$

### v2

@@ 4.1 General Convergence Analysis
Using the same restrictions on local and global learning rates placed by the FedAvg convergence analysis in (Karimireddy et al.,, 2020), the aforementioned bounded gradient variance, bounded gradient dissimilarity, and L-smoothness assumptions ensure that our method converges to $\mathbf{w}^{*}$ faster than FedAvg. Similar convergence rate arguments for other FL methods with appropriately modified aggregation can be shown, as discussed in Sec. 3.
@@ 4.2 Effectiveness of FedBaF with Diverse Client Data
This section shows the impact of integrating a foundation model close to the optimal weights on the learning process and convergence behavior in non-IID settings.
@@
Aggregating the updated models from clients according to Alg. 1 Lines 8-10 forms the global model $\mathbf{w}^{\prime}_{t}$. By the triangle inequality, we get
$\|\mathbf{w}^{\prime}_{t}-\mathbf{w}^{*}\|\leq\delta_{t}.$
@@
Let $\mathbf{w}^{*}$ be a (bounded) local minimum of the global objective function in (1). Consider an FL algorithm that converges to $\mathbf{w}^{*}$ and let $\mathbf{w}^{\prime}_{t}$ be its global model. Consider FedBaF based on the same FL algorithm (with appropriately modified client updates and Lines 8-10 in Alg. 1) and let $\mathbf{w}_{t}$ be the FedBaF global model. FedBaF’s global model error has an upper bound of $\|\mathbf{w}_{t}-\mathbf{w}^{*}\|\leq\frac{\delta_{t}+\alpha_{t}\tau_{t}\gamma}{1+\alpha_{t}\tau_{t}}<\delta_{t}$.
@@ 5 Experimental Evaluations
In this section, we conduct a detailed evaluation of FedBaF’s performance on both local and global test datasets, comparing its performance to the no foundation model and weight initialization baseline algorithms for training FL models.
@@
More detailed testing results are provided in Appendix C, including the 1) computational efficiency of FedBaF and 2) additional evaluations using foundation models of varying quality trained on different amounts of data and real-world foundation model weights that are publicly available online.
@@ 5.1 Experimental Setup
Our experiments with popular image classification tasks encompass experiments using the CIFAR-10 and Rome Weather Image (Vaz,, 2021) datasets. We use Pre-ResNet and Vision Transformer (Dosovitskiy et al.,, 2021) architectures as Vision Transformers are popular foundation model architectures known for achieving remarkable performance (Zhou et al.,, 2024) for ImageNet challenges (Dosovitskiy et al.,, 2021). We train the foundation models on the Tiny ImageNet-200 and Weather Image (Xiao,, 2021) datasets. Our evaluations consider both IID and non-IID settings (see Appendix A). To demonstrate FedBaF’s generalizability to other tasks, we also evaluate it on a next-word prediction task using a Transformer language model pre-trained on the WikiText-2 dataset and tested on the Penn Treebank dataset.
@@
Details of other deep neural network architectures employed in our experiments and additional training specifics are provided in Tables 1 and 2 in Appendix A.
@@ 5.2 Experimental Results
Figures 3, 4, and 5 display the extensive test accuracy/perplexity evaluation results for FedBaF and our two baselines (no foundation model and weight initialization). We evaluate the three methods using both FedAvg and FedProx (Li et al.,, 2020) as the base FL training algorithms. We incorporate FedBaF into FedProx by modifying Line 8 and the ClientUpdate routine in Algorithm 1, including both IID and non-IID settings, with one non-adversarial scenario and four adversarial scenarios. We use FedProx’s aggregation step: $\mathbf{w}^{\prime}_{t+1}\leftarrow\sum_{k\in S_{t}}\frac{n_{k}}{\sum_{k\in S_{t}}n_{k}}\mathbf{w}_{t+1}^{k}-\mu(\mathbf{w}_{t+1}^{k}-\mathbf{w}_{t})$. Here, $\mu$ represents the regularization term that controls the trade-off between the local and global objectives. FedBaF then incorporates the foundation model weights as in Lines 9-11 of Algorithm 1.
5.2.1 Testing FedBaF Performance
@@
Figures 3 and 4 respectively show test accuracies of all three methods using FedAvg and FedProx for both Pre-ResNet and the Vision Transformer. In comparison to Pre-ResNet trained with no foundation model, FedBaF improved global model accuracy by 1.3% for FedAvg and 1.6% for FedProx in IID scenarios, and by 21.8% for FedAvg and 22.6% for FedProx in non-IID scenarios. For the Vision Transformer, FedBaF improved global performance relative to no foundation model by 10.8% for FedAvg and 0.0% for FedProx in IID settings and by 37.5% for FedAvg and 15.8% for FedProx in non-IID settings. We observe that both FedAvg and FedProx benefit from FedBaF’s inclusion of the foundation model, with particular benefits in more challenging scenarios with non-IID client data. Incorporating the foundation model mitigates slower FL convergence caused by non-IID data.
@@
In the next-word prediction task using a Transformer, FedBaF significantly outperformed training with no foundation model, reducing perplexity by 76.0% with FedAvg, whereas weight initialization yielded a 67.8% decrease relative to training without a foundation model with FedAvg.
Collectively, these findings indicate negligible differences between the test accuracies attained by FedBaF and those achieved with weight initialization. Simultaneously, FedBaF achieves privacy advantages as, unlike weight initialization, it does not reveal the foundation model weights to FL clients. Moreover, FedBaF showed better testing performance than the weight initialization or no foundation model methods in 10 out of the 14 experiment settings.
In Appendix C.3, we provide additional examples of FedBaF achieving privacy and test performance improvements. They include comparisons with FedAdam (Reddi et al.,, 2021), a state-of-the-art FL method leveraging pre-trained models.
5.2.2 FedBaF’s Robustness to Attacks
@@
FedBaF remains effective in maintaining robustness when faced with misclassification attacks and shows a more modest performance degradation in comparison to both baseline methods.
@@
We evaluate the robustness of FedBaF and our two baselines in the presence of adversarial clients. As the proportion of attacking clients ($\zeta$ in Figures 34, and 5) increases, test accuracy declines in all cases. This matches our intuition since neither FedBaF nor the two baselines are designed to perfectly defend against these attacks, which become more effective as more clients act as attackers.
@@
The test accuracy for the no foundation model baseline method drops significantly when 50% of the clients are attackers and the attack intensity $\lambda=5$. For Pre-ResNet under such attacks, the global performance drop, when compared to the case with no attackers, is 11.8% for FedAvg and 11.2% for FedProx in IID scenarios and 38.2% for FedAvg and 37.4% for FedProx in non-IID scenarios. With the Vision Transformer, these accuracy drops are 32.4% for FedAvg and 34.2% for FedProx in IID and 21.9% for FedAvg and 25.0% for FedProx in non-IID scenarios. Thus, training with no foundation model results in vulnerability to attacks in both IID and non-IID settings. This matches our expectations since it has no built-in defenses.
@@
In contrast, FedBaF experiences a much more modest performance decline under attack, demonstrating its robustness. For Pre-ResNet, FedBaF’s global performance decreases by 6.0% for FedAvg and 6.4% for FedProx in IID settings and by 16.5% for FedAvg and 16.3% for FedProx in non-IID settings, where the decrease is again measured for the most intense attack ($\zeta=50\%,\lambda=5$) relative to no attack. These accuracy drops are less than half of those experienced by the no foundation model method. With the Vision Transformer, FedBaF’s global performance decreases by only 4.9% for FedAvg and FedProx in IID settings and by 25% for both FedAvg and FedProx in non-IID settings.
@@
Finally, we comment that the results also show that the weight initialization method, which naïvely incorporates the foundation model and reveals the foundation model weights to clients, does confer some robustness to attacks. Weight initialization shows a performance decrease of 6.4% for FedAvg and 6.9% for FedProx in IID settings and 17.5% for FedAvg and 17.2% for FedProx in non-IID settings globally for Pre-ResNet. With the Vision Transformer, the decreases are 14.6% for both FedAvg and FedProx in IID settings and 18.4% for both FedAvg and FedProx in non-IID settings globally.
@@
These findings demonstrate that, similar to weight initialization, FedBaF offers considerable attack robustness compared to training with no foundation model. In 8 out of 12 cases, FedBaF suffers the minimal loss in performance when compared with the other two baselines, indicating its superior effectiveness in maintaining robustness under adversarial conditions.
@@ 6 Conclusion
This section includes supplementary experimental results, analyzing the effect of varying foundation model quality on FedBaF’s performance and comparisons using the official pre-trained foundation model. Computational complexity is also compared to scenarios without foundation models and weight initialization. Additionally, we include further experiments employing FedAdam (Reddi et al.,, 2021).
@@ C.2 Using the Official Pre-Trained Foundation Model
C.3 Generalized Applicability of FedBaF with FedAdam
To further assess the generalized applicability of FedBaF, we conducted additional experiments using the FedAdam Reddi et al., (2021) algorithm. Specifically, we evaluated its performance on image classification tasks with CIFAR-10 and Rome Weather Image datasets using Pre-ResNet and Vision Transformer, as well as on a language modeling task with a Transformer model pre-trained on the WikiText-2 dataset. We followed the same experimental setup, attack scenarios, and hyperparameters described in Section 5 and Appendix A.
For FedAdam, we set the global aggregation update learning rate to 0.01 for CIFAR-10 and Rome Weather Image datasets and 0.1 for the WikiText-2 dataset. Additionally, we configured the momentum parameters as $\beta_{1}=0.9$ and $\beta_{2}=0.99$ for all experiments.
Table 7: Image classification test accuracy results for Pre-ResNet using the no foundation, weight initialization, and FedBaF methods employing FedAdam (best of 3 trials).
Table 8: Image classification test accuracy results for Vision Transformer using no foundation, weight initialization, and FedBaF methods employing FedAdam (best of 3 trials).
Table 9: Next-word prediction perplexity results for Transformer models using no foundation, weight initialization, and FedBaF methods employing FedAdam (best of 3 trials). Lower perplexity is better.
Tables 7, 8, and 9 present the experimental results obtained using FedAdam across different datasets, model architectures, and attack scenarios. The results demonstrate that FedBaF remains effective across various federated optimization frameworks, reinforcing its adaptability in federated learning.
Among the 70 tested cases, FedBaF achieved similar or superior performance in 68 cases when compared to weight initialization approaches, outperforming cases without foundation models. This trend aligns with the findings in Section 5, further highlighting FedBaF’s robustness across different optimization strategies and experimental settings.
These findings confirm the reliability of FedBaF in federated learning, ensuring its applicability to both vision and language tasks under diverse conditions.
C.4 Computational Complexity
@@
Compared to the no foundation model cases, FedBaF requires significantly fewer computations across both IID and non-IID scenarios. Specifically, for Pre-ResNet IID and non-IID cases, computations are reduced by 88.1-91.4% and 89.7-92.4% for FedAvg, and by 88.2-91.4% and 89.8-92.4% for FedProx, respectively. Similarly, for Vision Transformer, IID and non-IID scenarios see a reduction of 96.9-97.3% and 93.9-96.8% for FedAvg, and 96.5-97.0% and 94.5-97.1% for FedProx in computations. Therefore, these findings indicate that FedBaF’s computational demands are relatively minimal despite the foundation model being integrated into every training round.
@@ E.1.1 Assumptions
Foundation Model Alignment:
The foundation model shares the same architecture as the global model, guaranteeing seamless integration during aggregation.
@@ F.2 Proof of Proposition 2
where we set $\gamma=\|\mathbf{w}_{\text{pre}}-\mathbf{w}^{*}\|$. Since non-IID data can cause significant variance in local updates, we compare the derived bound to FedAvg, where the bound on the distance between $\mathbf{w}_{t}$ and $\mathbf{w}^{*}$ is $\delta_{t}$. By assumption, $\gamma\leq\delta_{t}$ for earlier rounds (small $t$). We get $\frac{\delta_{t}+\alpha_{t}\tau_{t}\gamma}{1+\alpha_{t}\tau_{t}}\leq\delta_{t}$, which is equivalent to
@@ G.1 Possibility of Extracting a Foundation Model
$\|\mathbf{w}_{t}-\mathbf{w}_{t}^{\prime}\|$ $=\left\|\frac{1}{1+{\alpha}_{t}\tau_{t}}(\mathbf{w}_{t}^{\prime}+{\alpha}_{t}\tau_{t}\mathbf{w}_{pre})-\mathbf{w}_{t}^{\prime}\right\|$
$=\frac{{\alpha}_{t}\tau_{t}}{1+{\alpha}_{t}\tau_{t}}\|\mathbf{w}_{t}^{\prime}-\mathbf{w}_{pre}\|$
$\|\mathbf{w}_{t}-\mathbf{w}_{pre}\|$ $=\left\|\frac{1}{1+{\alpha}_{t}\tau_{t}}\mathbf{w}_{t}^{\prime}+{\alpha}_{t}\tau_{t}\mathbf{w}_{pre}-\mathbf{w}_{pre}\right\|$
$=\frac{1}{1+{\alpha}_{t}\tau_{t}}\|\mathbf{w}_{t}^{\prime}-\mathbf{w}_{pre}\|$
At the onset of training, both distances are equivalent since we make a strategic choice for the weight $\alpha_{t}\tau_{0}$ to be approximately 2. This simplifies the initial update rule for the global model such that the initial global model weights, $\mathbf{w}_{0}$, are an unweighted average of the client’s updated model weights $\mathbf{w}_{0}^{\prime}$ and the foundation model weights $\mathbf{w}_{pre}$. As the training progresses, ${\alpha}_{t}\tau_{t}$ typically decays to less than 1. We deduce for $t>0$:
$\frac{{\alpha}_{t}\tau_{t}}{1+{\alpha}_{t}\tau_{t}}<\frac{1}{1+{\alpha}_{t}\tau_{t}}\implies\|\mathbf{w}_{t}-\mathbf{w}_{t}^{\prime}\|<\|\mathbf{w}_{t}-\mathbf{w}_{pre}\|$